# Test-time Correlation Alignment

### v1

@@ Appendix B Proof of Theoretical Statement
Here, we present Theorem 3.5 again for convenience. {mdframed}[backgroundcolor=gray!10, topline=false, bottomline=false, leftline=false, rightline=false, innertopmargin=1pt, innerbottommargin=1pt] Theorem 3.5 Let $h_{\theta}(\cdot)=g(f(\cdot))$ be an L-Lipschitz continuous hypothesis on $\mathcal{H}$. $D_{s}$ and $D_{t}$ represent the source and test data, respectively. Let $\Omega:=\bigcup_{x\in\mathbb{D}_{t}}\mathcal{B}(x,r^{*})$ as the set of source instances near the test data, we sample $k$ instances from $\Omega$ and $\mathbb{D}_{t}$ to obtain $[X_{s},Z_{s},P_{s}]$ and $[X_{t},Z_{t},P_{t}]$ by $h_{\theta}(\cdot)$, respectively. Per 3.2, 3.3, and 3.4, with a probability of at least $1-exp(-\frac{k^{2}}{c_{t}\mu^{-}\pi_{d_{I}}(r^{*}/L)^{d_{I}}}+\log k)$, we have
@@
where $\pi_{d_{I}}=\lambda(\mathcal{B}(0,1))$ is the volume of the $d_{I}$ dimension unit ball and $d_{I}$ is the dimension of input $x$. Furthermore, considering the source correlation $\Sigma_{s}=\mathbb{E}[\tilde{Z_{s}}^{T}\tilde{Z_{s}}]$ and the test correlation $\Sigma_{t}=\tilde{Z_{t}}^{T}\tilde{Z_{t}}$, where $\tilde{Z}_{s}$ and $\tilde{Z}_{t}$ are the centered matrices. With a probability of at least $\min(1-exp(-\frac{k^{2}}{c_{t}\mu^{-}\pi_{d_{I}}(r^{*}/L)^{d_{I}}}+\log k),1-\delta)$, the correlation distance $\lVert\Sigma_{s}-\Sigma_{t}\rVert$ is bounded by:
$\lVert\Sigma_{s}-\Sigma_{t}\rVert_{F}\leq 2\lVert Z_{s}\rVert_{F}(\frac{\lVert\hat{Y}_{t}-P_{t}\rVert_{F}+A}{\lVert J_{g}(Z_{s})\rVert_{F}})+(\frac{\lVert\hat{Y}_{t}-P_{t}\rVert_{F}+A}{\lVert J_{g}(Z_{s})\rVert_{F}})^{2}+B$ (23)
@@
Let $\mathbb{I}(x_{s}\in\mathcal{B}(x_{t},r))$ be an indicator function, where $\mathbb{I}(x_{s}\in\mathcal{B}(x_{t},r))$ is independent and identically distributed Bernoulli random variables, representing the probability $\mathbb{D}_{s}(x_{s}\in\mathcal{B}(x_{t},r))$. Let $S_{n}(x_{t})=\sum_{i=1}^{n_{s}}\mathbb{I}(x_{s}\in B(x_{t},r))$ denotes the number of source instances $x_{s}\in\mathbb{D}_{s}$ that fall within $\mathcal{B}(x_{t},r)$. Then, $S_{n}(x_{t})$ follows a Binomial distribution. Let $W\sim\text{Binomial}(n_{s},c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}})$. By applying Chernoff’s inequality, we obtain the probability that the number of source data points falling within $\mathcal{B}(x_{t},r)$ is less than $m$:
$P(S_{n}(x_{t})<m)\leq P(W<m)=P(W-E[W]<-m)$
$\leq\exp\left(-\frac{m^{2}}{2E[W]}\right)=\exp\left(-\frac{m^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$ (29)
@@
$P(\|x_{s}^{(m)}-x_{t}\|\leq r)$ $=P(S_{n}(x_{t})\geq m)\geq 1-\exp\left(-\frac{m^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$ (30)
For a fixed $x_{t}$, we assume that its nearest neighbor $x_{s}$ has the same label, and thus set $m=1$. By applying the union bound, the desired probability can be expressed as follows:
@@
$\geq 1-k\exp\left(-\frac{1}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$
$=1-\exp\left(-\frac{1}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$
Thus, with at least the probability $1-\exp\left(-\frac{1}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$ (which is a tighter upper bound compared to Theorem 3.5, strengthening our theoretical results.), the distance satisfies $\|dx_{s}\|\leq r\leq r_{t}$.
@@
$1-\exp\left(-\frac{1^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$
$\geq 1-\exp\left(-\frac{k^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}}+\log k\right)$ (33)
@@
Next, we prove Equation 23. The sampled covariance matrix is given by:
$\hat{\Sigma}_{S}$ $=Z_{s}^{T}Z_{s}$ (35)
$\Sigma_{t}$ $=(Z_{s}+dZ_{s})^{T}(Z_{s}+dZ_{s})$
$=Z_{s}^{T}Z_{s}+Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}$ (36)
The change in the covariance matrix is:
$\Sigma_{t}-\hat{\Sigma}_{S}$ $=Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}$ (37)
@@
$\|\Sigma_{t}-\hat{\Sigma}_{S}\|_{F}$
$\leq\|Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}\|_{F}$
$\leq 2\|Z_{s}\|_{F}\|dZ_{s}\|_{F}+\|dZ_{s}\|_{F}^{2}$ (38)
Since we cannot determine the true $P_{s}$ from Equation 34, we scale $\|P_{t}-P_{s}\|_{F}$ as follows:
@@
where $l$ is the true labels. Additionally, $\hat{\Sigma}_{S}$ is obtained from $k$ source domain instances and contains statistical error relative to the true covariance matrix $\Sigma_{S}=E[\hat{\Sigma}_{S}]$. By Hoeffding’s inequality, we have:
$P(\|\Sigma_{S}-E[\hat{\Sigma}_{S}]\|_{F}^{2}\geq\epsilon)$ $\leq 2\exp\left(-\frac{2k\epsilon}{d^{2}}\right)$ (40)
Let $2\exp\left(-\frac{2k\epsilon}{d^{2}}\right)=\sigma$, then:
$\epsilon$ $=\frac{-d^{2}\log(\frac{\sigma}{2})}{2k}$ (41)
With a probability of at least $1-\sigma$, we have:
$\|\Sigma_{S}-E[\hat{\Sigma}_{S}]\|_{F}$ $<\sqrt{\epsilon}=d\sqrt{\frac{\log(2/\delta)}{2k}}$ (42)
Finally, combining Equations 34, B.1 and 42, we derive the following proposition: with at least $1-\exp\left(-\frac{1^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)\geq 1-\exp\left(-\frac{k^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}}+\log k\right)$:
@@
where $\hat{Y}_{t}$ is the one-hot encoding of $P_{t}$, $A=\|o(kr^{*})\|_{F}+\epsilon(h(X_{t}))+\epsilon(h(X_{s}))$ represents the output generalization error, and $B=d\sqrt{\frac{\log(2/\delta)}{2k}}$ is the sampling error.
@@ D.2 Backbones
Both ResNet and ViT backbones are well-established in the literature and serve as strong candidates for evaluating domain adaptation techniques, with ResNet-18/50 being more computationally efficient and ViT-B/16 being particularly effective in capturing complex relationships across domains.
@@ D.3 Implementation Details
For hyper-parameter selection in Domain Generalization task, we first identify the optimal parameter set based on the highest accuracy achieved on the default domain (art paintings in PACS and art in OfficeHome). These parameters are then applied to other domains to assess their performance. Specifically, we conduct a search for the learning rate within the range {1e-7, 5e-7, 1e-6, 5e-6, 1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1}. For methods that include an entropy filter component (e.g., TSD), we explore the entropy filter hyperparameter in the set {1, 5, 10, 15, 20, 50, 100, 200, 300}. For the LinearTCA method, we optimized the number of pseudo-source instances $k$ within the range {5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 100, 200, 300}. For most datasets and backbones, smaller $k$ values generally yield satisfactory results. For datasets with a substantial number of images per class, it is advisable to experiment with larger $k$ values. For the LinearTCA+ method, we conducted an optimization of $k$ values on the basis of other top-performing test-time adaptation method and its parameter settings.
For the Image Corruption task, each Test-Time Adaptation (TTA) method continually adapts to 15 corruptions following the specified order: [Gaussian Noise, Shot Noise, Impulse Noise, Defocus Blur, Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic Transformation, Pixelate, JPEG Compression]. We experiment with each TTA method using learning rates from {1e-7, 5e-7, 1e-6, 5e-6, 1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1} and the entropy filter hyperparameter in the set {1, 5, 10, 15, 20, 50, 100, 200, 300}. The parameter range for $k$ in the LinearTCA/LinearTCA+ methods remains consistent with the selection in Domain Generalization. The top-performing test-time adaptation approach on the Image Corruption is selected as the base method for LinearTCA+. The best performance results obtained for each method are selected as the final experimental outcomes.
During the Test-Time Adaptation phase, both the Domain Generalization and Image Corruption tasks utilize specific batch sizes for different backbones. ResNet-18 and ResNet-50 use a batch size of 128, whereas the ViT-B/16 is configured with a batch size of 64.
@@ E.1 Comparison Results Details
Tables 5, 6, 7, 8, 9 and 10 provide the detailed results of our experimental results on Domain Generalization task, and Tables 11, 12, 13, 14, 15 and 16 offers a detailed overview of the outcomes from our Image Corruption task. These results demonstrate that our TCA method consistently outperforms other state-of-the-art TTA approaches across most domians and corruption types, effectively validating the TCA’s capability to robustly enhance accuracy performance during the test phase.
@@ A C P S
Table 5: Accuracy comparison of different TTA methods on PACS based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
@@
Table 6: Accuracy comparison of different TTA methods on PACS based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
@@
Table 7: Accuracy comparison of different TTA methods on PACS based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.

### v2

@@ Appendix B Proof of Theoretical Statement
[…]
@@
where $\pi_{d_{I}}=\lambda(\mathcal{B}(0,1))$ is the volume of the $d_{I}$ dimension unit ball and $d_{I}$ is the dimension of input $x$. Furthermore, considering the true source correlation $\Sigma_{s}=\mathbb{E}[\tilde{Z_{s}}^{T}\tilde{Z_{s}}]$ and the pseudo-source correlation $\hat{\Sigma}_{s}=\tilde{Z_{t}}^{T}\tilde{Z_{t}}$, where $\tilde{Z}_{s}$ and $\tilde{Z}_{t}$ are centered. With a probability of at least $\min(1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right),1-\delta)$, the correlation distance $\lVert\Sigma_{s}-\hat{\Sigma}_{s}\rVert$ is bounded by:
$\lVert\Sigma_{s}-\hat{\Sigma}_{s}\rVert_{F}\leq$
@@
Let $\mathbb{I}(x_{s}\in\mathcal{B}(x_{t},r))$ be an indicator function, where $\mathbb{I}(x_{s}\in\mathcal{B}(x_{t},r))$ is independent and identically distributed Bernoulli random variables, representing the probability $\mathbb{D}_{s}(x_{s}\in\mathcal{B}(x_{t},r))$. Let $S_{n}(x_{t})=\sum_{i=1}^{n_{s}}\mathbb{I}(x_{s}\in B(x_{t},r))$ denotes the number of source instances $x_{s}\in D_{s}$ that fall within $\mathcal{B}(x_{t},r)$. Then, $S_{n}(x_{t})$ follows a Binomial distribution. Let $W\sim\text{Binomial}(n_{s},c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}})$. By applying Chernoff’s inequality, we obtain the probability that the number of source data points falling within $\mathcal{B}(x_{t},r)$ is less than $m$:
$P(S_{n}(x_{t})<m)=P(W<m)\leq\exp\left(-\frac{(E[W]-m)^{2}}{2E[W]}\right)=\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-m)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$ (29)
@@
$P(\|x_{s}^{(m)}-x_{t}\|\leq r)$ $=P(S_{n}(x_{t})\geq m)\geq 1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-m)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$ (30)
For a fixed $x_{t}$, it suffices to find a single nearest neighbor $x_{s}$ that lies within the ball $B(x_{t},r)$, and thus we set $m=1$. By applying the union bound, the desired probability can be expressed as follows:
@@
$\geq 1-k\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}\right)$
$=1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$
Thus, with at least the probability $1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$, the distance satisfies $\|dx_{s}\|\leq r\leq r_{t}$.
@@
$1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$
@@
Next, we prove Equation 23. Let $\Sigma_{s}^{\prime}$ denote the correlation matrix computed from $k$ sampled source instances $Z_{s}$, and let $\hat{\Sigma}_{s}$ denote the pseudo-source correlation matrix computed from $k$ sampled test instances $Z_{t}$. These matrices are computed as follows:
$\Sigma_{s}^{\prime}$ $=Z_{s}^{T}Z_{s}$ (34)
$\hat{\Sigma}_{s}=Z_{t}^{T}Z_{t}=(Z_{s}+dZ_{s})^{T}(Z_{s}+dZ_{s})$ $=Z_{s}^{T}Z_{s}+Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}$ (35)
The change in the correlatione matrix is:
$\hat{\Sigma}_{s}-\Sigma_{s}^{\prime}$ $=Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}$ (36)
@@
$\|\hat{\Sigma}_{s}-\Sigma_{s}^{\prime}\|_{F}\leq\|Z_{s}^{T}dZ_{s}+(dZ_{s})^{T}Z_{s}+(dZ_{s})^{T}dZ_{s}\|_{F}\leq 2\|Z_{s}\|_{F}\|dZ_{s}\|_{F}+\|dZ_{s}\|_{F}^{2}$ (37)
Additionally, since $\Sigma_{s}^{\prime}$ is obtained from $k$ source domain instances and contains statistical error relative to the true covariance matrix $\Sigma_{s}=E[\Sigma_{s}^{\prime}]$. By Hoeffding’s inequality, we have:
$P(\|\Sigma_{s}^{\prime}-E[\Sigma_{s}^{\prime}]\|_{F}^{2}\geq\epsilon)$ $\leq 2\exp\left(-\frac{2k\epsilon}{d^{2}}\right)$ (38)
Here, $d$ denotes the range of $\Sigma_{s}^{\prime}$, which is set to 1. Let $2\exp\left(-\frac{2k\epsilon}{d^{2}}\right)=\sigma,$ then:
$\epsilon$ $=-\frac{\log(\frac{\sigma}{2})}{2k}$ (39)
With a probability of at least $1-\sigma$, we have:
$\|\Sigma_{s}^{\prime}-\Sigma_{s}\|_{F}$ $<\sqrt{\epsilon}=\sqrt{\frac{\log(2/\delta)}{2k}}$ (40)
By combining Equations 37 and 40, we obtain:
$\|\Sigma_{s}-\hat{\Sigma}_{s}\|_{F}\leq\|\Sigma_{s}-\Sigma_{s}^{\prime}\|_{F}+\|\Sigma_{s}^{\prime}-\hat{\Sigma}_{s}\|_{F}\leq\sqrt{\frac{\log(2/\delta)}{2k}}+2\|Z_{s}\|_{F}\|dZ_{s}\|_{F}+\|dZ_{s}\|_{F}^{2}$ (41)
We can further expand Equation 41 by applying Equation 33. However, since we cannot determine the true $P_{s}$ in Equation 33, we scale $\|P_{t}-P_{s}\|_{F}$ as follows:
@@
where $l$ is the true labels.
Finally, combining Equations 33, B.1 and 41, we derive the following proposition: with at least $min(1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right),1-\sigma)$:
@@
where $\hat{Y}_{t}$ is the one-hot encoding of $P_{t}$, $A=\|o(kr^{*})\|_{F}+\epsilon(h(X_{t}))+\epsilon(h(X_{s}))$ represents the output generalization error, and $B=\sqrt{\frac{\log(2/\delta)}{2k}}$ is the sampling error.
@@ D.1 Datasets
  • DomainNet: The DomainNet dataset is a large-scale dataset used in transfer learning, consisting of 6 domains: {clipart, infograph, painting, quickdraw, real, and sketch}. It consists of a total of 586,575 images, with each domain containing 345 classes.
@@
  • ImageNet-C: ImageNet-C is significantly larger compared to CIFAR10-C and CIFAR100-C. This dataset contains 1,281,167 training images and 50,000 test images, categorized into 1,000 classes. Like CIFAR10-C and CIFAR100-C, ImageNet-C also includes 15 types of corruptions.
@@ D.2 Backbones
  • CLIP: Contrastive Language- Image Pre-Training (CLIP), developed by OpenAI, is a cutting-edge multimodal model that bridges visual and textual domains through contrastive learning. CLIP employs dual encoders (ResNet/ViT for images and Transformer for text) to project both modalities into a shared semantic space, enabling zero-shot classification by matching image features with natural language prompts.
Both ResNet and ViT backbones are well-established in the literature and serve as strong candidates for evaluating domain adaptation techniques, with ResNet-18/50 being more computationally efficient and ViT-B/16 being particularly effective in capturing complex relationships across domains. In this work, the zero-shot classification model CLIP is also included as a backbone to validate the effectiveness of our proposed methods on closed-source foundation models.
@@ D.3 Implementation Details
Consistent with prior work (Wang et al., 2020; Niu et al., 2022, 2023; Nguyen et al., 2023; Yuan et al., 2024; Iwasawa & Matsuo, 2021; Wang et al., 2023; Zhang et al., 2023), hyperparameter tuning in our experimental setup is conducted across datasets. Specifically, in the Domain Generalization task, we first identify the optimal parameter set based on the highest accuracy achieved on the default domain (art paintings in PACS, art in OfficeHome and clipart in DomainNet). These parameters are then applied to other domains to assess their performance. Specifically, we conduct a search for the learning rate within the range {1e-7, 5e-7, 1e-6, 5e-6, 1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1}. For methods that include an entropy filter component (e.g., TSD), we explore the entropy filter hyperparameter in the set {1, 5, 10, 15, 20, 50, 100, 200, 300}. For AdaNPC, we explore the hyperparameter $k$ (the number of nearest samples used for voting) over {5, 10, 15, 20, 30, 40, 50}. For the LinearTCA method, we optimized the number of pseudo-source instances $k$ within the range {5, 10, 15, 20, 25, 30, 35, 40, 45, 50, 100, 200, 300}. For most datasets and backbones, smaller $k$ values generally yield satisfactory results. For datasets with a substantial number of images per class, it is advisable to experiment with larger $k$ values. For the LinearTCA+ method, we conducted an optimization of $k$ values on the basis of other top-performing test-time adaptation method and its parameter settings.
For the Image Corruption task, we experiment with each TTA method using learning rates from {1e-7, 5e-7, 1e-6, 5e-6, 1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1} and the entropy filter hyperparameter in the set {1, 5, 10, 15, 20, 50, 100, 200, 300}. The parameter range for $k$ in AdaNPC, LinearTCA/LinearTCA+ remains consistent with their respective selections in Domain Generalization task. The top-performing test-time adaptation approach on the Image Corruption is selected as the base method for LinearTCA+. The best performance results obtained for each method are selected as the final experimental outcomes. For the pre-trained model on ImageNet-C dataset, we utilize the model provided by TorchVision.
During the Test-Time Adaptation phase, both the Domain Generalization and Image Corruption tasks utilize specific batch size for different backbones. ResNet-18 and ResNet-50 use a batch size of 128, whereas the ViT-B/16 is configured with a batch size of 64.
@@ E.1 Comparison Results Details
Tables 9, 10, 11, 12, 13, 14, 15, 16 and 17 provide the detailed results of our experimental results on Domain Generalization task, and Tables 18, 19, 20, 21, 22, 23, 24, 25 and 26 offers a detailed overview of the outcomes from our Image Corruption task. These results demonstrate that our TCA method consistently outperforms other state-of-the-art TTA approaches across most domians and corruption types, effectively validating the TCA’s capability to robustly enhance accuracy performance during the test phase.
@@ A C P S
Table 9: Accuracy comparison of different TTA methods on PACS dataset based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
@@
Table 10: Accuracy comparison of different TTA methods on PACS dataset based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
@@
Table 11: Accuracy comparison of different TTA methods on PACS dataset based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.